# StableMotion: Training Motion Cleanup Models with Unpaired Corrupted Data

### v1

@@ 6. Motion Cleanup
Once the generate-discriminate model has been trained, it can be used to clean raw mocap data by detecting corrupted segments and inpainting them with clean motions, while leaving the uncorrupted frames unchanged. As described in Section 5.2, the model can function as a motion quality evaluator $D({\mathbf{h}}|{\mathbf{m}})$. Given a raw mocap sequence ${\mathbf{m}}$ as input, the model evaluates each frame to determine if a frame exhibits artifacts. Frames that require cleaning are identified as ${\mathbb{A}}\leftarrow\{i\mid D({\mathbf{h}}^{i\in 1:N}|{\mathbf{m}})=\mathbf{1}\}$. The input for motion inpainting is then constructed by masking out the segments $({\mathbf{m}},{\mathbf{h}})^{i\in{\mathbb{A}}}$ and assigning QualVar as ${\mathbf{h}}\leftarrow\mathbf{0}$, which specifies high-quality motions for all frames. The model is then used to generate clean motion for all corrupted frames $G({\mathbf{m}}^{i\in{\mathbb{A}}}|{\mathbf{m}}^{i\in{\mathbb{B}}},\mathbf{0}^{1:N})$. In our experiments, we observe that the generate-discriminate model exhibits strong motion cleanup capabilities, but can lead to drastic changes in the content of the original motion in some scenarios due to the masked-inpainting process. Furthermore, the inherent stochasticity of diffusion models can occasionally produces suboptimal results. In this section, we describe a number of test-time techniques designed to address these limitation and further improve the consistency of our system.
Figure 3. To improve preservation of motion content, we propose an adaptive cleanup technique using soft motion quality evaluation and soft motion inpainting. A simple Monte Carlo method is used to estimate soft quality labels $\bar{{\mathbf{h}}}$ from the predicted QualVar ${\mathbf{h}}$, which approximates the severity of artifacts in each frame. Then, we design a soft inpainting strategy that inpaints corrupted frames by initializing the denoising process for each frame ${\mathbf{m}}^{i}$ at different diffusion steps ${{\mathbf{t}}^{i}_{\text{soft}}}$. When correcting a frame with subtle artifacts (i.e. low $\bar{{\mathbf{h}}}^{i}$), the inpainting process is initialized at a later denoising step ${\mathbf{t}}^{i}_{\text{soft}}\textless T$, allowing more information from the original frame to be retained.
@@ 6.1. Adaptive Cleanup
To identify the level of corruption in each frame, we approximate the soft quality label $\bar{{\mathbf{h}}}^{i}$ by taking multiple samples of $\hat{{\mathbf{h}}}^{i}$ for each frame, and then computing the expected value over the samples. This approach provides a soft estimate of the level of corruption in each frame, instead of relying directly on the value of a single sample. This approximation assumes that severely corrupted frames are easier to identify, resulting in higher expected values of the indicator variable, while frames with more subtle artifacts tend to yield lower expected values.
Once the level of corruption for each frame has been estimated, we use the diffusion-denoising process to generate motions starting from an intermediate diffusion step ${\mathbf{t}}^{i}_{\text{soft}}$ from each frame rather than the maximum step $T$. The initial denoising step ${\mathbf{t}}^{i}_{\text{soft}}$ is determined based on the soft quality labels $\bar{{\mathbf{h}}}^{i}$. In practice, we observe that starting from later denoising steps (e.g., ${\mathbf{t}}_{\text{soft}}<\frac{T}{2}$) often fails to fully remove artifacts. To address this, we design a soft inpainting schedule that encourages inpainting from earlier diffusion steps, while still retaining useful information from the original corrupted frames. The initial denoising step is determined according to:
(3) $\small{\mathbf{t}}^{i}_{\text{soft}}\begin{cases}T\sin{\frac{\pi}{2}\min\left(1,2\ \mathbb{E}(\hat{{\mathbf{h}}}^{i})-1+\tau\right)},&\text{if }\mathbb{E}(\hat{{\mathbf{h}}}^{i})\geq\tau\\ 0,&\mathrm{otherwise}\\ \end{cases},$
where $\mathbb{E}(\hat{{\mathbf{h}}}^{i})$ denotes the Monte-Carlo estimate of the soft quality label $\bar{{\mathbf{h}}}^{i}$ for the $i$-th frame, and $\tau=0.5$ serves as the threshold to determine whether a frame should be inpainted. A visualization of the soft inpainting schedule is illustrated in Figure 4. With this schedule, soft inpainting is performed when $\mathbb{E}(\hat{{\mathbf{h}}}^{i})$ lies between $0.5$ and $0.75$, while for $\mathbb{E}(\hat{{\mathbf{h}}}^{i})>0.75$, the full denoising process will be used starting from diffusion step $T$. This adaptive cleanup technique allows the model to adaptively adjust the modification to each frame based on the severity of the artifacts, enabling it to better preserve the content of the original frames.
@@ 6.2. Quality-Aware Ensemble
To further improve the consistency and performance of our system, we leverage the sample diversity of diffusion models through ensembling, where results from multiple diverse models are combined to improve generalization (Kotu and Deshpande, 2019). However, instead of generating samples from different models, we propose to ensemble different predictions from a single model. An overview of this ensemble process is illustrated in Figure 5. We first execute the motion cleanup process multiple times, consisting of motion evaluation and motion inpainting, and collect the candidate output motions. Since each DDPM inference process starts with independent Gaussian noise, the outcomes are diverse and independent. However, naively averaging the results is unsuitable for motion data, as averaging across diverse modes can introduce artifacts and distort the natural dynamics of the motion. Instead, the generate-discriminate model enables selection among the candidates based on the estimated motion quality. The motion evaluation procedure $D({\mathbf{h}}|{\mathbf{m}})$ is applied to each candidate motion to estimate the quality of each candidate. The overall quality score of a motion sequence is computed as the sum of ${\mathbf{h}}$ at each frame and provides an automatic criterion for selecting the best quality candidate. In our experiments, we generate 5 candidate results for ensembling. Since multiple candidate motions can be generated in parallel using batch inference, this approach does not significantly increase the overall wall-clock time of the motion cleanup process. More detailed ablation experiments that evaluate the effectiveness of our quality-aware ensembling approach is available in Section E.2.
@@ 7. Model Representations
In this section, we describe the implementation decisions for several crucial components of the system.
Motion Representation. The commonly used relative trajectory representation can lead to error accumulation when reconstructing global motion (Holden et al., 2016). To localize edits and preserve uncorrupted frames during cleanup, we adopt a global motion representation inspired by prior work on human mesh recovery and motion cleanup (Zhang et al., 2024; Hwang et al., 2025), where each frame is represented by the following features: global root translation, ground-projected facing direction, local joint rotations and positions, global foot positions and velocities, and global root velocity. Here, “global” features are defined relative to the first frame, which is canonicalized to ensure consistent initial orientation. This results in a 244-dimensional feature vector ${\mathbf{m}}^{i}$ per frame for SoccerMocap, further augmented with a 1D quality indicator ${\mathbf{h}}^{i}$.
Network Architecture Our model is implemented using a diffusion transformer (DiT) (Peebles and Xie, 2023). To better capture relative temporal correlations between frames, we adopt rotary positional embeddings (RoPE) in the attention layers (Su et al., 2024; Shen et al., 2024; Evans et al., 2024), instead of absolute positional embeddings. More details are available in Appendix B.
@@ 8. Real-world Experiments: SoccerMocap
We start by evaluating the effectiveness of our framework on SoccerMocap, a real-world mocap dataset from a prominent game studio.
@@
SoccerMocap is a proprietary, large-scale soccer dataset captured using optical and marker-free motion capture systems, including full-pitch performances. It contains roughly 245 hours of motion data recorded at 50 fps, and exhibits various artifacts introduced by in-the-wild capture conditions, such as large camera-to-subject distance, occlusion, system failures, and equipment vibrations. Since the majority of corrupted segments are much shorter than 5 seconds, we set the maximum context length of our neural network to 5 seconds. At test time, long motion clips can be cleaned by applying our model with a sliding window.
@@
QualVar for training are annotated by a joint effort of artists and heuristic algorithms. Frames containing artifacts are first annotated by artists, accounting for approximately 1% of the frames. These annotations focus on challenging artifacts, such as body self-penetration and abnormal pose errors, which can be difficult to detect using heuristic algorithms. We refer to these Human-Annotated artifacts as HAA in further discussion. To expand the annotations, we apply heuristic artifact detection algorithms designed to identify foot-skating, flipping, motion pops, and frozen frames. This process results in approximately 39% of the frames being identified as corrupted. These heuristics are designed to achieve high recall at the expense of lower precision compared to human annotations, ensuring that nearly all true corrupted frames are identified. The dataset is split into a 240-hour training set and a 5-hour testing set.
Results. Cleanup models are trained using StableMotion directly on the SoccerMocap training set, and then applied to clean the testing set. Examples of motions from SoccerMocap before and after cleanup by the model are illustrated in Figure 1 and Figure 7. For the kicking motion shown in Figure 1, it can take 30–60 minutes for a professional animator to fix the body self-penetration, whereas our model detects and corrects it automatically with reasonably natural and consistent results. For artifacts such as frozen frames and motion pops shown in Figure 7 (A) and Figure 6, which could take hours for animators to fix, our model is able to automatically remove these artifacts. The model has been used to automatically process newly captured soccer mocap data from similar systems. The kicking motion shown in Figure 1 serves as an example from a recent capture session.
Table 1. Performance of StableMotion framework on SoccerMocap. A proprietary motion artifact detection algorithm is applied to evaluate motion quality before and after motion cleanup. FS Dist measures the severity of foot skating. Pops Rate quantifies the frequency of jittery motions and motion pops. Frozen Rate captures the frequency of over-smoothed transitions and frozen frames. R@HAA denotes the Recall of model-predicted QualVar for Human-Annotated Artifacts, with more details available in Appendix A. This score primarily reflects the model’s sensitivity in identifying artifacts that are difficult for automatic algorithms to detect, such as body self-penetration and abnormal poses.
Method FS Dist $\downarrow$ Pops Rate $\downarrow$ Frozen Rate $\downarrow$ Jittering $\downarrow$ R@HAA $\uparrow$
Input 2.39 0.41 % 23.86 % 8.94 -
CondMDI-B 3.25 18.21 % 4.64% 22.09 -
StableMotion 2.14 0.13 % 4.50 % 8.68 81.54 %
w/ Clean-Only 2.96 0.23 % 4.43 % 14.67 72.74 %
w/o QualVar 3.76 1.91 % 6.43 % 9.85 -
Since previous methods require either clean or paired corrupted-to-clean data to train cleanup models, most cannot be directly trained on SoccerMocap, a raw mocap dataset that contains artifacts. For a fair comparison, we build several baseline models, including one trained only on filtered clean data (w/ Clean-Only), and the other trained on mixed-quality data but without QualVar (w/o QualVar). We also implement a state-of-the-art motion inpainting model, CondMDI (Cohan et al., 2024), trained on mixed-quality data, denoted as CondMDI-B. For fairness, all baseline models use the same motion representation and use the corrupted frame labels predicted by our StableMotion model. To quantitatively evaluate performance, we use an internal motion artifact detection algorithm to estimate artifact severity before and after cleanup. As shown in Table 1, FS Dist, Pops Rate, Frozen Rate, and Jittering measure the levels of foot skating, motion pops, frozen frames, and jittery behaviors, respectively. R@HAA reflects the model’s ability to accurately identify artifacts that typically require human annotation and are challenging for heuristic algorithms to detect. We refer the reader to Appendix A for more details about these metrics. The results from CondMDI-B and the “w/o QualVar” baseline show that directly training models on raw data can exhibit, or even amplify, motion artifacts—potentially due to the influence of artifacts, such as jittery or popping, in the training data. Simply filtering out corrupted frames and training the motion generation component of our model on the resulting clean but fragmented snippets can degrade motion quality, possibly due to the loss of temporal coherence. In contrast, StableMotion effectively trains a motion cleanup model directly on raw data using the quality indicators, preserving the continuity and structure of the motion sequences. The model is able to fix a wide range of motion artifacts introduced by real-world mocap systems and exhibits more effective performance than all baselines. The high R@HAA further highlights the model’s ability to accurately detect subtle artifacts that typically require human annotation.
@@ 9. Benchmark Experiments: BrokenAMASS
To further evaluate the effectiveness of training a cleanup model from corrupted data in a controlled setting, we create BrokenAMASS by introducing artifacts into the AMASS dataset to simulate real-world raw mocap data that contains artifacts without clean paired motions (Mahmood et al., 2019). This dataset contains over 40 hours of motion data recorded at 20 fps. We design a collection of corruption functions, including random-joint perturbation, over-acceleration, drifting, and over-smoothing, to simulate real-world motion artifacts, such as jittering, foot sliding, and other unnatural movements. Detailed implementations of these corruption functions are provided in the appendix. The default BrokenAMASS dataset contains approximately 23% corrupted frames. Models are trained on the training set and evaluated on the testing set. The dataset is split into training, test, and validation sets with a 0.80 : 0.15 : 0.05 ratio.
Metrics. BrokenAMASS serves as a benchmark for evaluating content preservation by enabling direct comparison between the cleaned-up motions produced by different models and the corresponding clean ground-truth motions from AMASS (Mahmood et al., 2019). Following prior work, we also adopt several heuristic metrics that estimate the severity of common motion artifacts, such as foot skating and jittering. The evaluation metrics include: FS Dist for foot sliding distance (Ling et al., 2020), Jitter from the first derivative of acceleration (Shen et al., 2024), and Accel for acceleration error compared to clean motion (Zhang et al., 2024). GMPJPE assesses joint position errors in global frames. M2M Score and M2M R@3 quantify motion content semantic similarity using the TMR model (Petrovich et al., 2023). GMPJPE and M2M provide complementary metrics to evaluate content preservation and high-level semantic consistency respectively.
Table 2. Quantitative comparison on BrokenAMASS. The performance of each method is averaged across three different random seeds, and reported along with its standard deviation. Boldface indicates the best method, and an underscore denotes the second best. $\rightarrow$” indicates that a value closer to the ground truth (when available) represents better performance. Methods marked with “*” use the ground truth quality labels to determine which frames need cleanup.
@@
RoHM ${4.43}^{\pm{0.11}}$ $\underline{{2.02}}^{\pm{0.15}}$ $\underline{{1.00}}^{\pm{0.03}}$ $\underline{{7.56}}^{\pm{0.63}}$ $\mathbf{{0.940}}^{\pm{0.002}}$ $\underline{{81.2}}^{\pm{0.54}}$
StableMotion $\mathbf{{3.70}}^{\pm{0.01}}$ $\mathbf{{1.14}}^{\pm{0.04}}$ $\mathbf{{0.60}}^{\pm{0.01}}$ $\mathbf{{2.42}}^{\pm{0.04}}$ $\underline{{0.938}}^{\pm{0.003}}$ $\mathbf{{84.5}}^{\pm{0.77}}$
Comparisons. We compare the performance of our method against strong baseline methods from prior work, each trained under their intended dataset conditions using either clean or paired corrupted-to-clean motion data: (1) CondMDI, a conditional motion in-betweening diffusion model trained on clean data (Cohan et al., 2024); and (2) RoHM, a non-inpainting diffusion-based motion restoration method trained with paired corrupted-to-clean data (Zhang et al., 2024). To ensure a fair comparison, we train each baseline on the type of data that best matches its original design and supervision assumptions. Since existing methods are not designed to learn directly from raw, corrupted motion data, we additionally construct a variant of CondMDI trained on mixed-quality unpaired data, which we denote as CondMDI-B. As shown in Table 2, CondMDI trained on corrupted data performs worse than the version trained on clean data. This performance gap indicates that the CondMDI framework relies on training with clean data to be effective, limiting its applicability in scenarios where obtaining clean datasets is challenging. In contrast, our model is trained only on an unpaired corrupted dataset, but still outperforms state-of-the-art models for motion cleanup that are trained with cleaned motion data. The motions corrected by our model exhibit even fewer foot skating artifacts compared to the clean motions from the original AMASS dataset. Moreover, our method is able to better preserve the content of the original motions by applying the cleanup process only to the corrupted frames. This is particularly crucial for highly dynamic locomotion behaviors, as illustrated in Figure 8. Compared to RoHM, the non-inpainting model trained on paired data, our method demonstrates stronger performance for fixing artifacts, achieving a 16% reduction in foot skating distance and a 40% reduction in acceleration error. At the same time, it achieves comparable performance in content preservation, with comparable semantic preservation performance and a 68% reduction in global joint position errors.

### v2

@@ 6. Motion Cleanup
Once the generate-discriminate model has been trained, it can be used to clean raw mocap data by identifying corrupted frames and inpainting them with clean motions, while leaving the uncorrupted frames unchanged. As described in Section 5.2, the model can function as a motion quality evaluator $D({\mathbf{h}}|{\mathbf{m}})$. Given a raw mocap sequence $\mathbf{m}$} as input, the model evaluates each frame to determine if a frame exhibits artifacts. Frames that require cleaning are identified as ${\mathbb{A}}\leftarrow\{i\mid D({\mathbf{h}}^{i\in 1:N}|{\mathbf{m}})=\mathbf{1}\}$. The input for motion inpainting is then constructed by masking out the segments $({\mathbf{m}},{\mathbf{h}})^{i\in{\mathbb{A}}}$ and assigning QualVar as ${\mathbf{h}}\leftarrow\mathbf{0}$, which specifies high-quality motions for all frames. The model is then used to generate clean motion for all corrupted frames $G({\mathbf{m}}^{i\in{\mathbb{A}}}|{\mathbf{m}}^{i\in{\mathbb{B}}},\mathbf{0}^{1:N})$. In our experiments, we observe that the generate-discriminate model exhibits strong motion cleanup capabilities, but can lead to drastic changes in the content of the original motion in some scenarios due to the masked-inpainting process. Furthermore, the inherent stochasticity of diffusion models can occasionally produces suboptimal results. In this section, we describe a number of test-time techniques designed to address these limitation and further improve the consistency of our system.
Figure 3. To improve preservation of motion content, we propose an adaptive cleanup technique using soft motion quality evaluation and soft motion inpainting. A simple Monte Carlo method is used to estimate soft quality labels $\bar{{\mathbf{h}}}$ from the predicted QualVar $\mathbf{h}$}, which approximates the severity of artifacts in each frame. Then, we design a soft inpainting strategy that inpaints corrupted frames by initializing the denoising process for each frame ${\mathbf{m}}^{i}$ at different diffusion steps ${\mathbf{t}}^{i}_{\text{soft}}$}. When correcting frames with subtle artifacts (i.e. low $\bar{{\mathbf{h}}}^{i}$), the inpainting process is initialized at later denoising steps ${\mathbf{t}}^{i}_{\text{soft}}\textless T$, thereby retaining more information from the original frames.
@@ 6.1. Adaptive Cleanup
To identify the level of corruption in each frame, we approximate a soft quality label $\bar{{\mathbf{h}}}^{i}$ by taking multiple samples of $\hat{{\mathbf{h}}}^{i}$ for each frame, and then computing the expected value over the samples. This approach provides a soft estimate of the level of corruption in each frame, instead of relying directly on the value of a single sample. This approximation assumes that severely corrupted frames are more consistently identified as being corrupted, resulting in higher expected values of the indicator variable, while frames with more subtle artifacts tend to yield lower expected values.
Once the level of corruption for each frame has been estimated, we use the diffusion-denoising process to generate motions starting from different intermediate diffusion steps ${\mathbf{t}}^{i}_{\text{soft}}$ at each frame rather than the maximum step $T$. The initial denoising step ${\mathbf{t}}^{i}_{\text{soft}}$ is determined based on the soft quality labels $\bar{{\mathbf{h}}}^{i}$. In practice, we observe that starting from later denoising steps (e.g., ${\mathbf{t}}_{\text{soft}}<\frac{T}{2}$) often fails to fully remove artifacts. To address this, we design a soft inpainting schedule that encourages inpainting from earlier diffusion steps, while still retaining useful information from the original corrupted frames. The initial denoising step is determined according to:
(3) $\small{\mathbf{t}}^{i}_{\text{soft}}=\begin{cases}T\sin{\frac{\pi}{2}\min\left(1,2\bar{{\mathbf{h}}}^{i}-1+\tau\right)},&\text{if }\bar{{\mathbf{h}}}^{i}\geq\tau\\ 0,&\mathrm{otherwise}\\ \end{cases},$
where $\bar{{\mathbf{h}}}^{i}\approx\mathbb{E}[\hat{{\mathbf{h}}}^{i}]$ denotes the Monte-Carlo estimate of the soft quality label for the $i$-th frame, and $\tau=0.5$ is the threshold to determine whether a frame should be inpainted. A visualization of the soft inpainting schedule is illustrated in Figure 4. With this schedule, soft inpainting is performed when $\bar{{\mathbf{h}}}^{i}$ lies between $0.5$ and $0.75$, while for $\bar{{\mathbf{h}}}^{i}>0.75$, the full denoising process will be applied starting from diffusion step $T$. This adaptive cleanup technique allows the model to adaptively adjust the modification to each frame based on the severity of the artifacts, enabling it to better preserve the content of the original frames.
@@ 6.2. Quality-Aware Ensemble
To further improve the consistency and performance of our system, we leverage the sample diversity of diffusion models through ensembling, where results from multiple diverse models are combined to improve generalization (Kotu and Deshpande, 2019). However, instead of generating samples from different models, we propose to ensemble different predictions from a single model. An overview of this ensemble process is illustrated in Figure 5. We first execute the motion cleanup process multiple times, consisting of motion evaluation and motion inpainting, and collect the candidate output motions. Since each DDPM inference process starts with independent Gaussian noise, the outcomes are diverse and independent. However, naively averaging the results is unsuitable for motion data, as averaging across diverse modes can introduce artifacts and distort the natural dynamics of the motion. Instead, the generate-discriminate model enables selection among the candidates based on the estimated motion quality. The motion evaluation procedure $D({\mathbf{h}}|{\mathbf{m}})$ is applied to each candidate motion to estimate the quality of each candidate. The overall quality score of a motion sequence is computed as the sum of $\mathbf{h}$} at each frame and provides an automatic criterion for selecting the best quality candidate. In our experiments, we generate 5 candidate motions for ensembling. Since multiple motions can be generated in parallel using batch inference, this approach does not significantly increase the overall wall-clock time of the motion cleanup process. More detailed ablation experiments that evaluate the effectiveness of our quality-aware ensembling approach is available in Section E.2.
@@ 7. Model Representations
In this section, we describe the implementation details for several crucial components of the system.
Motion Representation.
The commonly used relative trajectory representation can lead to error accumulation when reconstructing global motions (Holden et al., 2016; Tevet et al., 2023). To localize edits and preserve uncorrupted frames during cleanup, we adopt a global motion representation inspired by prior work on human mesh recovery and motion cleanup (Zhang et al., 2024; Hwang et al., 2025). With this global representation, each frame is represented by features consisting of: global root translation, ground-projected facing direction, local joint rotations and positions, global foot positions and velocities, and global root velocity. Here, “global” features are defined relative to the first frame, which is canonicalized to ensure consistent initial orientation. This results in a 244-dimensional feature vector ${\mathbf{m}}^{i}$ per frame for SoccerMocap, which is further augmented with a 1D quality indicator ${\mathbf{h}}^{i}$.
Network Architecture.
Our model is implemented using a diffusion transformer (DiT) (Peebles and Xie, 2023). Adaptive normalization is used to condition the model on the diffusion step and a class label (Chen et al., 2024), which indicates whether the motion or QualVar is observed during inpainting. To better capture relative temporal correlations between frames, we adopt rotary positional embeddings (RoPE) in the attention layers (Su et al., 2024; Shen et al., 2024). More architectural details are available in Appendix B.
@@ 8. Real-world Experiments: SoccerMocap
We start by evaluating the effectiveness of our framework on SoccerMocap, a real-world mocap dataset from a prominent game studio. SoccerMocap is a proprietary, large-scale soccer dataset captured using optical and marker-free motion capture systems in full-field environments. It contains approximately 245 hours of motion data recorded at 50 fps, and exhibits artifacts introduced by in-the-wild capture conditions, such as large camera-to-subject distances, occlusions, system failures, and equipment vibrations. Since the majority of corrupted segments are much shorter than 5 seconds, the maximum context length of our model is set to 5 seconds. Once trained, long motion clips can be cleaned by applying our model with a sliding window.
@@
The quality indicator variables used for training are annotated by a joint effort of artists and heuristic algorithms. Frames containing artifacts are first annotated by artists, accounting for approximately 1% of the frames. These annotations focus on challenging artifacts, such as body self-penetration and abnormal pose errors, which can be difficult to detect using heuristic algorithms. We refer to these human-annotated artifacts as HAA in the following discussion. To expand the annotations, we apply heuristic artifact detection algorithms designed to identify foot-skating, flipping, motion pops, and frozen frames. This process results in approximately 39% of the frames being identified as corrupted. These heuristics are designed to achieve high recall at the expense of lower precision compared to human annotations, ensuring that nearly all corrupted frames are correctly identified. The dataset is split into a 240-hour training set and a 5-hour testing set.
@@
Results
StableMotion models are trained directly on the SoccerMocap training set, and then applied to clean the testing set. Examples of motions from SoccerMocap before and after cleanup by the model are illustrated in Figure 1 and Figure 7. For the kicking motion shown in Figure 1, it can take 30–60 minutes for a professional animator to fix the body self-penetrations, whereas our model detects and corrects it automatically with reasonably natural and consistent results. Our model is able to automatically remove artifacts such as frozen frames and motion pops, shown in Figure 7 (A) and Figure 6, which can take hours for animators to fix. The model has been used to automatically process newly captured soccer mocap data from similar systems. The kicking motion shown in Figure 1 is an example from a recent capture session.
Table 1. Performance of the StableMotion on SoccerMocap. A proprietary artifact detection algorithm is used to evaluate motion quality before and after cleanup. FS Dist measures the severity of foot skating. Pops Rate quantifies the frequency of jittery transitions and motion pops. Frozen Rate captures the frequency of over-smoothed transitions and frozen frames. StableMotion provides a highly effective method for training motion cleanup models directly on raw, corrupted datasets in real-world scenarios.
Method FS Dist $\downarrow$ Pops Rate $\downarrow$ Frozen Rate $\downarrow$ Jittering $\downarrow$
Input 2.39 0.41 % 23.86 % 8.94
XClean 2.86 0.35 % 23.07 % 15.18
ConvAE-B 9.54 47.46 % 5.07 % 104.44
CondMDI-B 3.25 18.21 % 4.64% 22.09
w/o QualVar 3.76 1.91 % 6.43 % 9.85
[…]
Since previous methods require either clean or paired corrupted-to-clean data to train cleanup models, most cannot be directly trained on SoccerMocap, a raw mocap dataset with pervasive artifacts. To evaluate StableMotion’s effective utilization of raw motion data, we compare StableMotion against two ablations of our method: one trained only on filtered clean data (w/ Filtered), and the other trained on mixed-quality data without using QualVar (w/o QualVar). To compare StableMotion with a traditional non-learning based cleanup technique, we implement an example-based cleanup method (XClean) (Hsu et al., 2004; Arikan and Forsyth, 2002), commonly used in game studio pipelines. We also compare our method with a state-of-the-art motion inpainting model, CondMDI (Cohan et al., 2024), trained on raw, mixed-quality data without text conditions, which we refer to as CondMDI-B. To compare with inpainting-based methods developed prior to the rise of diffusion models (Harvey and Pal, 2018; Harvey et al., 2020; Hernandez et al., 2019; Kaufmann et al., 2020), we constructed a convolutional autoencoder baseline following Kaufmann et al. (2020), denoted as ConvAE-B, and reimplement it using a more modern network architecture. All baseline models use the same motion representation. Since not all baseline methods have the ability to identify corrupted frames, the corrupted frame labels predicted by our model is used for all baseline methods. To quantitatively evaluate performance, we use a proprietary motion artifact detection algorithm to estimate artifact severity before and after cleanup. As shown in Table 1, FS Dist, Pops Rate, Frozen Rate, and Jittering measure the levels of foot skating, motion pops, frozen frames, and jittery behaviors, respectively. We refer the reader to Appendix A for more details about these metrics.
@@ Results
ConvAE, CondMDI-B, and the “w/o QualVar” baseline show that directly training models on raw data can exhibit, or even amplify, motion artifacts—potentially due to the influence of artifacts, such as jittery or popping, in the training data. The results from “w/ Filtered” further show that simply filtering out corrupted frames and training the motion generation component of our model on the resulting clean but fragmented snippets can degrade motion quality, possibly due to the loss of temporal coherency. In contrast, StableMotion effectively trains a motion cleanup model directly on raw data using the quality indicators, preserving the continuity and structure of the motion sequences. The model is able to correct a wide range of motion artifacts introduced by real-world mocap systems and exhibits more effective performance than the baselines. To evaluate our model’s accuracy in identifying artifacts, we compare the predicted quality indicators with the quality labels from human annotators. Our model accurately detect subtle artifacts that typically require manual annotation, achieving an HAA recall rate of 81.54%.
@@ 9. Benchmark Experiments: BrokenAMASS
To evaluate the effectiveness of StableMotion for training cleanup models from corrupted data in a controlled setting, we create BrokenAMASS by introducing artifacts into the AMASS dataset to simulate real-world raw mocap data that contains artifacts without clean paired motions (Mahmood et al., 2019). This dataset contains over 40 hours of motion data from the original AMASS dataset, downsampled to recorded at 20 fps. The dataset covers diverse motion types, including locomotion, transitions, sports, dance, yoga, and martial arts (Punnakkal et al., 2021). We design a collection of corruption functions, including random-joint perturbation, over-acceleration, drifting, and over-smoothing, to simulate real-world motion artifacts, such as jittering, foot sliding, and other unnatural movements. Detailed implementations of these corruption functions are provided in the Appendix. The default BrokenAMASS dataset contains approximately 23% corrupted frames. The dataset is split into training, test, and validation sets with a 0.80 : 0.15 : 0.05 ratio.
Metrics.
BrokenAMASS serves as a controlled benchmark for evaluating content preservation by enabling direct comparison between the cleaned-up motions produced by different models and the corresponding clean ground-truth motions from AMASS (Mahmood et al., 2019). Following prior work, we also adopt several heuristic metrics that estimate the severity of common motion artifacts, such as foot skating and jittering. The evaluation metrics include: FS Dist for foot sliding distance (Ling et al., 2020), Jitter from the first derivative of acceleration (Shen et al., 2024), and Accel for acceleration error compared to clean motion (Zhang et al., 2024). GMPJPE assesses joint position errors in global frames. M2M Score and M2M R@3 quantify motion content semantic similarity using the TMR model (Petrovich et al., 2023). GMPJPE and M2M provide complementary metrics to evaluate content preservation and high-level semantic consistency respectively.
Table 2. Quantitative comparison on BrokenAMASS. Results are averaged over three random seeds and reported with standard deviations. The best performing model is highlighted, and underlines denote the second-best. $\rightarrow$” means that values closer to the ground truth are better. Methods marked with “*” use ground-truth quality labels to identify frames for cleanup. Our method effectively removes motion artifacts while also preserving the original motion content.
@@
[…]
@@
Comparisons.
We compare the performance of our method against strong baseline methods from prior work, each trained under their intended dataset conditions using either clean or paired corrupted-to-clean motion data: (1) CondMDI, a conditional motion in-betweening diffusion model trained on clean data (Cohan et al., 2024); and (2) RoHM, a non-inpainting diffusion-based motion restoration method trained with paired corrupted-to-clean data (Zhang et al., 2024). To ensure a fair comparison, we train each baseline on the type of data that best matches its original design. Since existing methods are not designed to be trained directly on raw, corrupted motion data, we construct both a diffusion-based and a non-diffusion-based motion in-betweening model trained on mixed-quality unpaired data, which we denote as CondMDI-B and ConvAE-B (Kaufmann et al., 2020). As shown in Table 2, CondMDI trained on corrupted data performs worse than the version trained purely on clean data. This performance gap indicates that the CondMDI framework relies on training with clean data to be effective, limiting its applicability in scenarios where obtaining clean datasets is challenging. In contrast, our model is trained only on unpaired corrupted data, but still outperforms state-of-the-art cleanup models that are trained with cleaned motion data. The motions corrected by our model exhibit even fewer foot skating artifacts compared to the clean motions from the original AMASS dataset. Moreover, our method is able to better preserve the content of the original motions by applying the cleanup process only to the corrupted frames. This is particularly crucial for highly dynamic locomotion behaviors, as illustrated in Figure 8. Compared to RoHM, the non-inpainting model trained on paired data, our method demonstrates stronger performance for fixing artifacts, achieving a 16% reduction in foot skating distance and a 40% reduction in acceleration error. StableMotion also achieves comparable performance in content preservation, with similar semantic preservation performance and a 68% reduction in global joint position errors.
Qualitative Results.